1.A survey from teenage research unlimited found that 80% of teenage consumers received their spending money from part-time jobs. If five teenagers are selected at random, find the probability that at least three of them will have part-time jobs


2.PMEA associate market research found that 40% of Nigerians don’t think that having a college education is important to succeed in the business world, if a random sample of five Nigerians are selected, find these probabilities


A.That exactly 2 people will agree with that statement


B.At most 3 people will agree with that statement


C.At least 2 people will agree with that statement


D.Fewer than 3 people will agree with that statement


That none of the people will agree with that statement

1
Expert's answer
2021-11-17T14:41:30-0500

1. Let "X=" the number of teenagers who will have part-time jobs: "X\sim Bin (n, p)."

Given "n=5, p=0.8, q=1-p=0.2"


"P(X\geq3)=P(X=3)+P(X=4)+P(X=5)"

"=\dbinom{5}{3}(0.8)^3(0.2)^{5-3}+\dbinom{5}{4}(0.8)^4(0.2)^{5-4}"

"+\dbinom{5}{5}(0.8)^5(0.2)^{5-5}=0.2048+0.4096+0.32768"

"=0.94208"

2. Let "X=" the number of people who will agree with that statement: "X\sim Bin (n, p)."

Given "n=5, p=0.4, q=1-p=0.6"

A.


"P(X=2)=\dbinom{5}{2}(0.4)^2(0.6)^{5-2}=0.3456"

B.


"P(X\leq3)=1-P(X=4)-P(X=5)"

"=1-\dbinom{5}{4}(0.4)^4(0.6)^{5-4}+\dbinom{5}{5}(0.4)^5(0.6)^{5-5}"

"=1-0.0768-0.01024=0.91296"

C.


"P(X\geq2)=1-P(X=0)-P(X=1)"

"=1-\dbinom{5}{0}(0.4)^0(0.6)^{5-0}+\dbinom{5}{1}(0.4)^1(0.6)^{5-1}"

"=1-0.07776-0.2592=0.66304"

D.



"P(X<3)=P(X=0)+P(X=1)+P(X=2)"

"=\dbinom{5}{0}(0.4)^0(0.6)^{5-0}+\dbinom{5}{1}(0.4)^1(0.6)^{5-1}"

"+\dbinom{5}{2}(0.4)^2(0.6)^{5-2}"

"=0.07776+0.2592+0.3456=0.68256"

E.


"P(X=0)=\dbinom{5}{0}(0.4)^0(0.6)^{5-0}=0.07776"